Hi,
I hope you can help. Many thanks for your time. Hugh
Make k the subject of each of the following two equations. Show each step
of your working.
Question A
2m=9t/5k−7
Question B
5k = 3t− 3k/t

Answer by Theo(13342)

i suppose you mean to solve for k.
let's take a look:

first problem:
start with:
2M = (9T / 5K) - 7
add 7 to both sides of this equation to get:
2M + 7 = 9T / 5K
multiply both sides of this equation by K to get:
K * (2M + 7) = 9T / 5
divide both sides of this equation by (2M + 7) to get:
K = 9T / (5 * (2M + 7))
you can simplify this further to get:
K = 9T / (10M + 35)
that's your answer. assuming your original equation is as shown above.

second problem:
start with:
5K = 3T - (3K / T)
multiply both sides of this equation by T to get?
5KT = 3T^2 - 3K
add 3K to both sides of this equation to get:
5KT + 3K = 3T^2
factor out the K on the left side of this equation to get:
K * (5T + 3) = 3T^2
divide both sides of this equation by (5T + 3) to get:
K = 3T^2 / (5T + 3)
that's your answer, assuming your original equation is as shown above.
I don't believe it can be simplified any further.